at what rate to the nearest hundreth of a percent, would you need to invest 13,000 compounded continuously to reach 21,000 in 7 years?
Answer by stanbon(75887)

at what rate to the nearest hundreth of a percent, would you need to invest 13,000 compounded continuously to reach 21,000 in 7 years?
----
A(t) = P*e^(rt)
21000 = 13000*e^(7r)
---
e^(7r) = 21/13
---
Take the natural log of both sides to get:
7r = ln(21/13)
7r = 0.4796
rate = 0.0685 = 6.85%
==========================
Cheers,
Stan H.